OpenStudy (anonymous):

Make: "r" the subject of A = pi r^2 (that is "pi" times "r^2)

OpenStudy (angela210793):

A = pi r^2 : pi both sides A/pi=r^2(sqr both sides r=sqr(A/pi)

OpenStudy (anonymous):

Looks good, thank you.

OpenStudy (angela210793):

uw :)